Jimmy has a brokerage account and buys on the margin, which resulted in an interest expense of $36,400 during the year. Income generated through the brokerage account was as follows:

Municipal interest $72,800
Taxable dividends and interest 364,000

If required, round any division to two decimal places and use in subsequent computation. Round your final answer to the nearest dollar.

How much investment interest can Jimmy deduct?
Answer #1

Investment interest can Jimmy deduct = [$364,000 / ($72,800 + $364,000)] × $36,400 = $30,333.33
Answer #2

SOLUTION :



Income from Jimmy’s brokerage account :



Taxable dividends and interest = 364000 ($)


Non-taxable interest = 72800 ($)


Total interest = Taxable part + Non-taxable part = 364000 + 72800 = 436800  ($)


Taxable portion of the total = 364000 / 436800 = 5 / 6 



Interest expense of Jimmy :



Jimmy has total interest expense  = 36400 ($)


So,


that portion of interest expense which correspond to the taxable portion of same amount of interest income of $36400 = 36400 * (5/6) =  30333.33 ($) is deductible.


So, Jimmy can deduct  $30333.33 ($) (ANSWER). 

answered by: Tulsiram Garg